OpenStudy (anonymous):

Solve x^2 - 3x = -8 Lemme see if I get this right... x^2 - 3x = -8 (x-1.5)^2 = -8 -1 (x-1.5)^2 = -9 x - 1.5 = +or- sqrt9i x-1.5 = 3sqrt2?

OpenStudy (anonymous):

I am horribly wrong. These are my choices. http://prntscr.com/92694d

OpenStudy (tkhunny):

You're very close. Give it another go. x^2 - 3x = -8 x^2 - 3x + (3/2)^2 = -8 + (3/2)^2 (x - (3/2))^2 = -8 + (9/4) = -23/4 x - (3/2) = +/- sqrt(-23)/2 x = (3/2) +/- sqrt(-23)/2 x = (3/2) +/- i*sqrt(23)/2

OpenStudy (anonymous):

I don't get what to do.

OpenStudy (anonymous):

And the end of your reply is not an answer choice.

OpenStudy (anonymous):

@tk

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@pooja195 @kohai

OpenStudy (anonymous):

@Kainui @Hero

OpenStudy (tkhunny):

Your first step needs improvement. You can't just declare it a perfect square, you must make it one. Compare your first step to my fist two steps. Where did you get "-1"?

OpenStudy (anonymous):

I got -1.5 by dividing 3x by 2.

OpenStudy (tkhunny):

You had 3 You divided by 2 to get 3/2. (3/2)^2 = 9/4 Where did you get -1?

OpenStudy (anonymous):

Oh I had a typo there. I meant -1.5 and I guess I stuck with it by accident.

OpenStudy (anonymous):

I wanted to do it to both sides :/

OpenStudy (anonymous):

What do I do?

OpenStudy (tkhunny):

It's never negative. You must ADD 9/4 to both sides.

OpenStudy (anonymous):

Can ya walk me through the rest? I'm really bad at Algebra II.

OpenStudy (anonymous):

Hello?

OpenStudy (tkhunny):

Already did. It's in my first post.

OpenStudy (anonymous):

Is it 3 + or - isqrt23/2?

OpenStudy (tkhunny):

(3 + or - isqrt23)/2 or 3/2 + or - isqrt23/2? Take your pick.

OpenStudy (tkhunny):

Learning just a little LaTeX goes a VERY long way. \(x = \dfrac{-3\pm i\sqrt{23}}{2}\) or \(x = \dfrac{3}{2}\pm\dfrac{i\sqrt{23}}{2}\)